Meeting Minutes 10th July 2013
Minutes from 10/07 (no meeting on 3rd.)
Attendees :- Stevie Conlon, Dave Berry, Peter Jackson, Neil Grant, Michael Sun,
Apologies : Tim Gray, Richard Good
Update on Current Tasks
- DR testing strategy determined and test arranged.
- TEST server work nearing completion
- LIVE server build in progress.
AOB
The remaining tasks for DEV Tech determined to be
- DR request of ITI ,
- Test the SOA dehydration database message tidy scripts
- Complete server builds
- Additional days (2) required for standby database creation as part of new infrastructure.
Load Testing and Review of monitoring tools have been dropped.
Actions
8. TAD to be updated ongoing with completion expected w/c 06/05 . To be Signed Off at Dev tech weekly Meeting , 09/05. TAD to contain policy of SOA DB and log archiving limit of 60 days. Required for build sign off.
Action Neil :
Update on 5th :- TAD to complete when all known tasks resolved on TEST.
Update on the 26th . TAD to be circulated Thursday 27th for review.
Update on the 11th TAD to be reviewed at the Dev Tech meeting on the 12th July.
16. Outstanding JIRA to be closed and completed by Dev Tech Team. Preferred for build sign off.
Action : Richard/Michael – complete for Development Team.
Action Neil : 2 outstanding JIRAs with Dev Tech
Update on 26th. Ongoing to be reviewed as part of the build review.
Update on the 10th. Solutions to outstanding JIRAs to be implemented and tested.
17. Arrange build review and sign off
Action : Stevie
Update for 12th May. Awaiting TAD. To arrange w/c 17th.
Update for the 26th . Build Review arranged for the 3rd July. PPIR by email for 4th July.
Update for the 3rd. PBBR not signed and TAD to be completed and JIRA solutions implemented. New milestones planned as follows.
Revised milestones as follows
PPBR 15 July (was 3rd July )
PPIR 22 July (was 4th July)
Complete LIVE builds 23rd July (was 8th July)
DSOR/Closure 26th July (was 19th July)
22.Peter to initiate DR test with ITI involvement. As the service is not resilient, this will involve bringing down the current test server and bringing it up elsewhere. This will require around 1 days downtime for the test server..
Action Peter
Update for the 10th. DR test schedule with ITI for the 18th July on the SOA TESTserver.
25. Follow up outstanding Oracle call with Oracle contact
Action Stevie/Michael
Update for 10th July. Query has been passed to Oracle contact. The current solution is acceptable but Oracle may be able to suggest a better way in the future.
26. Peter to test SOA Dehydration Database Housekeeping scripts on Friday 28th.
Action Peter
Update for 10th July. Peter prepped for testing on Friday 28th. Peter has been testing the scripts on the Dev database.
Update for 17th.
Closed Actions
21. Determine Tech Man Involvement. Arrange discusson with Dev Tech and Tech Man to allow Heather L to determine involvement as no running applications will be on the servers at present.
Action Stevie, Dev Tech,Tech Man.
Update required for 26th. See action 24. Closed
23. Report potential issue to Oracle (as described above)
Action Michael.
Update required for 25th June. Complete .Call raised with Oracle
24,Meet with Tech Man ( Heather Larnach) to verify their requirements from the project.
Action : Stevie
Update. Stevie met with Heather who will allocate resource to review the TAD and servers with Dev Tech. Closed